OpenStudy (anonymous):

0.6 kilograms = how many grams

Parth (parthkohli):

0.6 kilograms = 0.6 * 1 kilogram = 0.6 * 1000 grams = ? grams

OpenStudy (anonymous):

multiply by 1000

Parth (parthkohli):

What is 0.6 * 1000?

OpenStudy (anonymous):

and change it to grams.

OpenStudy (anonymous):

6000

OpenStudy (anonymous):

you are off by one zero.

OpenStudy (anonymous):

600

OpenStudy (anonymous):

Yes!! Be careful when multiply decimal values

OpenStudy (anonymous):

ok